# Automated modal analysis of entanglement with bipartite self-configuring optics

### v1

@@ Abstract
Entanglement is a unique feature of quantum mechanics. In coupled systems of light and matter, entanglement manifests itself in the linear superposition of multipartite quantum states (e.g., parametrized by the multiple spatial, spectral, or temporal degrees of freedom of a light field). In bipartite systems, the Schmidt decomposition provides a modal decomposition of the entanglement structure over independent, separable states. Although ubiquitous as a tool to describe and measure entanglement, there exists no general method to decompose a bipartite quantum state onto its Schmidt modes. Here, we propose a method that relies on bipartite self-configuring optics that automatically “learns” the Schmidt decomposition of an arbitrary pure quantum state. Our method is agnostic to the degrees of freedom over which quantum entanglement is distributed and can reconstruct the Schmidt modes and values by variational optimization of the network’s output powers or coincidences. We illustrate our method with numerical examples of spectral entanglement analysis for biphotons generated via spontaneous parametric down conversion and provide experimental guidelines for its realization, including the influence of losses and impurities. Our method provides a versatile and scalable way of analyzing entanglement in bipartite integrated quantum photonic systems.
@@
A successful implementation of quantum computing and simulation protocols entails an accurate description and characterization of entanglement. In pure bipartite quantum systems, measures of entanglement are provided by the Schmidt number and the von Neumann entropy horodecki2009quantum . Generally, the Schmidt decomposition of a pure bipartite state provides a convenient modal description of entanglement (as a linear superposition of separable and orthogonal “Schmidt modes”) fedorov2014schmidt ; sharapova2015schmidt . When the Hilbert space of photon pairs is low-dimensional, the quantum state and its underlying entanglement can be unveiled with a finite set of projective measurements toninelli2019concepts . However, when the Hilbert space becomes large and even continuous, as in the case of spatial and spectral correlations just2013transverse ; sharapova2015schmidt ; graffitti2020direct ; morrison2022frequency ; shukhin2024two , methods to experimentally measure the Schmidt modes are more challenging to implement. Such methods have been implemented for photon pairs with a priori knowledge of the mode shape straupe2011angular . Despite the ubiquity and usefulness of Schmidt modes in describing multimode entanglement, there is no general method to experimentally determine the Schmidt decomposition of a pure biphoton state.
@@
Here, we propose a general method to automatically analyze entanglement in bipartite quantum optical systems using self-configuring optics. Our approach relies on bipartite self-configuring networks (BSCN) composed of MZI meshes that impart unitary transformations to the two subspaces of the bipartite system. The average power or coincidence counts at the output of the BSCN are sequentially optimized, which results in the network having “learned” the Schmidt decomposition of an input state. Then, the Schmidt number can be directly measured at the output (as the number of ports with non-zero average power/coincidence), and the Schmidt modes of the system can be deduced directly from the network’s settings. Our approach thus automatically discovers the Schmidt modes without presuming any prior knowledge about these modes.
We illustrate our method with numerical examples of photon pairs generated via SPDC in various phase-matching settings. We show that our method can automatically measure Schmidt modes of photon pairs while being robust to single photon losses. Since our method allows us to automatically learn a modal decomposition of entanglement in the system, we also discuss potential applications in modal shaping of separable light and generation of states with controlled entanglement. Our method paves the way to full characterization, processing, and generation of entanglement in quantum optical states in a robust, scalable, and tunable manner. We envision that our approach could form a building block in integrated quantum optical architectures for computing and simulation where entanglement measurement and control plays an essential role.
@@
Figure 2: Sequential optimization methods for entanglement analysis in self-configuring optics. a-d. Sequential power optimization. a. Power optimization sequence: from $1$ to $N_{A}$ (on outputs of network $A$), then from $1$ to $N_{B}$ (on outputs of network $B$). b. Random matrix input example. c. Resulting eigenvector reconstruction from power optimization. d. Resulting simulated square roots of power outputs (meas. power) and singular values (sing. values) at the outputs of network $A$ (red) and $B$ (blue). e-h. Sequential coincidence counts optimization. e. Coincidence counts optimization sequence: from $1$ to $\text{min}(N_{A},N_{B})$ (on pairs of outputs ports from networks $A$ and $B$). f. Random matrix input example. g. Resulting eigenvector reconstruction from coincidence optimization. h. Resulting simulated square roots of coincidence counts (meas. coinc.) and singular values (sing. values).
@@
If single photons in either subspace are lost (e.g., absorbed or scattered), the output mixes with the vacuum state, which might result in errors in the average power measured. We show in the SM, Section S3 that, when modeling detection or scattering losses as coupling each output waveguide to a reservoir, the influence of losses can be calibrated out by measuring the average power transmission at each output port. Large losses may however result in reducing the number of counts on the detector and slow down the training process. Other MZI mesh architectures may also alleviate the influence of losses and imperfections burgwal2017using ; miller2015perfect . This intrinsic robustness to single photon losses provides further motivation to investigate the performance of our method in experimentally-relevant settings.
@@
Figure 4: Schmidt mode reconstruction and entangled light processing with bipartite self-configuring optics. a. Left: First four Schmidt modes reconstructed from BSCN parameters in the SPDC example of Figure 3d,g.. Right: BSCN architecture for reference, where each self-configuring layer is highlighted in a different color. The self-configuring layers used to reconstruct the Schmidt modes on the left are represented by the number and color of gray squares as insets. b. Once trained, BSCN can be used to generate separable light with of a given modal shape by blocking all output ports except the two ports corresponding to that mode index (in networks $A$ and $B$) and feeding them as input to networks applying the inverse transformation. c. If trained on a highly degenerate input state (such that two singular values are approximately equal), BSCN can also be used to generate supermode Bell states (by blocking all output ports except the four ports corresponding to degenerate singular values in each network). An additional phase shifter can be added on one of the output ports to control the relative phase between the two terms of the pair.
@@
In the absence of losses and spectral mixing by the gratings, the resulting input state to the BSCN is:
@@
A distinctive feature of our method is that the Schmidt modes (including their phases) are directly encoded into the parameters of the network after it has converged. Specifically, the $k$-th Schmidt mode is encoded in the network parameters (e.g., the phase shifts for various waveguide segments of MZIs) of layers $1,\ldots,k$ of both networks $A$ and $B$. We can therefore reconstruct the Schmidt modes (of the SPDC example from Figure 3(d)) as shown in Figure 4(a). The self-configuring layers used for the reconstruction of each mode are indicated with dashed square boxes in each panel.
@@ Discussion
Our methods could be applied to entanglement analysis in other quantum systems, such as temporal entanglement in atom-photon coupling bogdanov2006schmidt and macroscopic quantum states of light, such as bright squeezed vacuum sharapova2015schmidt ; sharapova2018bright . In both cases, multimodal entanglement over temporal, spatial, or spectral degrees of freedom can be described on the Schmidt basis which reveals the fundamental mechanisms of entanglement generation. Many of these systems, including photon pairs from SPDC, rely on several control parameters of light or matter, which can be dynamically optimized to shape or maximize entanglement. Since the BSCN automatically learns the Schmidt decomposition of the input state, it could also serve as a form of feedback on the system’s control parameters to learn their influence of the entanglement structure of the input state. In particular, well established techniques for shaping photon pairs u2006generation ; dosseva2016shaping ; meyer2017limits ; graffitti2020direct ; morrison2022frequency ; hurvitz2023frequency ; shukhin2024two are often sensitive to experimental parameters such as beam alignment, poling fabrication, and even temperature. Our proposal could thus allow for robustness and tunability even in the presence of experimental imperfections.
The automatic learning of the input state’s Schmidt decomposition entails a pure input state of the network |$\ket{\psi_{\text{in}}}$. Several sources of noise and imperfections may result in impurity of the quantum state. First, stochastic losses through the network yield a mixture of pure states with different photon numbers. With the above-mentioned loss model, we show that post-selection on a given number state (e.g., by performing coincidence or single photon counts) mitigates the influence of this impurity on the performance of our algorithm (see SM, Section S5). Second, the input state might be mixed with other (random) states, resulting in a statistical mixture. In general, this would result in the emergence of cross-coincidence counts between output ports |$\ket{x_{j}}_{A}$ and |$\ket{x_{k}}_{B}$ where $j\neq k$ after diagonalization of the reduced density matrices performed by the BSCN (see SM, Section S5). Practically, this means that the BSCN might also be used to “detect” impurities in the input quantum state in this more general setting.
@@
I Competing interests
@@
II Data and code availability statement
III Acknowledgements
The authors would like to thank Carson Valdez, Annie Kroo, Anna Miller, and Olav Solgaard for stimulating conversations. C. R.-C. is supported by a Stanford Science Fellowship. A.K. is supported by the VATAT-Quantum fellowship by the Israel Council for Higher Education; the Urbanek-Chodorow postdoctoral fellowship by the Department of Applied Physics at Stanford University; the Zuckerman STEM leadership postdoctoral program; and the Viterbi fellowship by the Technion. S. F. and D.A.B. M. acknowledge support by the Air Force Office of Scientific Research (AFOSR, grant FA9550-21-1-0312). D.A.B.M. also acknowledges support by the Air Force Office of Scientific Research (AFOSR, grant FA9550-23-1-0307).

### v2

@@ Abstract
Entanglement is a unique feature of quantum mechanics. In coupled systems of light and matter, entanglement manifests itself in the linear superposition of multipartite quantum states (e.g., parametrized by the multiple spatial, spectral, or temporal degrees of freedom of a light field). In bipartite systems, the Schmidt decomposition provides a modal decomposition of the entanglement structure over independent, separable states. Although ubiquitous as a mathematical tool to describe and measure entanglement, there exists no general efficient experimental method to decompose a bipartite quantum state onto its Schmidt modes. Here, we propose a method that relies on bipartite self-configuring optics that automatically “learns” the Schmidt decomposition of an arbitrary pure quantum state. Our method is agnostic to the degrees of freedom over which quantum entanglement is distributed and can reconstruct the Schmidt modes and values by variational optimization of the network’s output powers or coincidences. We illustrate our method with numerical examples of spectral entanglement analysis for biphotons generated via spontaneous parametric down conversion and provide experimental guidelines for its realization, including the influence of losses and impurities. Our method provides a versatile and scalable way of analyzing entanglement in bipartite integrated quantum photonic systems.
@@
A successful implementation of quantum computing and simulation protocols entails an accurate description and characterization of entanglement. In pure bipartite quantum systems, measures of entanglement are provided by the Schmidt number and the von Neumann entropy horodecki2009quantum . Generally, the Schmidt decomposition of a pure bipartite state provides a convenient modal description of entanglement (as a linear superposition of separable and orthogonal “Schmidt modes”) fedorov2014schmidt ; sharapova2015schmidt . Performing the Schmidt decomposition becomes progressively challenging as the Hilbert space becomes larger and even continuous, as in the case of encoding quantum entanglement in the spatial and spectral correlations of photon pairs just2013transverse ; sharapova2015schmidt ; graffitti2020direct ; morrison2022frequency ; shukhin2024two .
The problem of efficient Schmidt decomposition is an active field of study amooei2024efficient ; bhattacharjee2022measurement ; averchenko2020reconstructing ; kulkarni2018angular ; kulkarni2022classical , important for applications such as high-dimensional quantum communication amooei2024efficient ; dixon2012quantum ; luo2019quantum ; sit2017high and quantum imaging asban2019quantum ; fuenzalida2022resolution . Assumptions on the spatial and spectral shape and symmetries of the driving pump field and nonlinear crystal in SPDC can approximate the Schmidt decomposition analytically kulkarni2018angular ; kulkarni2022classical ; averchenko2020reconstructing ; sharapova2015schmidt ; sharapova2018bright ; straupe2011angular ; law2004analysis ; jha2011partial ; van2006effect ; miatto2012spatial and reduce the scaling of the number of required measurements amooei2024efficient ; bhattacharjee2022measurement . However, to date there is no efficient method for measuring the Schmidt decomposition in the most general setting of high-dimensional entangled photon states.
@@
Here, we propose a general method to automatically analyze entanglement in bipartite quantum optical systems using self-configuring optics. Our approach relies on bipartite self-configuring networks (BSCN) composed of MZI meshes that impart unitary transformations to the two subspaces of the bipartite system. The average power or coincidence counts at the output of the BSCN are sequentially optimized, which results in the network having “learned” the Schmidt decomposition of an input state. Then, the Schmidt number can be directly measured at the output (as the number of ports with non-zero average power/coincidence), and the Schmidt modes of the system can be deduced directly from the network’s settings. Our approach thus automatically discovers the Schmidt modes without presuming any prior knowledge about these modes. Furthermore, our architecture sequentially learns the most important Schmidt modes first (corresponding to the largest Schmidt values). In the practically important case where only a few Schmidt modes dominate, our method offers a favorable linear scaling with the Hilbert space dimension, both in required physical hardware and number of measurements, compared to the quadratic scaling if one directly measures the entire density matrix in order to perform a Schmidt decomposition.
We illustrate our method with numerical examples of photon pairs generated via SPDC in various phase-matching settings. We show that our method can automatically measure Schmidt modes of photon pairs while being robust to single photon losses. Since our method allows us to automatically learn a modal decomposition of entanglement in the system, we also discuss potential applications in modal shaping of separable light, generation of states with controlled entanglement, and in entanglement distribution protocols for quantum communication through scattering media. Our method paves the way to full characterization, processing, generation and distribution of entanglement in quantum optical states in a robust, scalable, and tunable manner. We envision that our approach could form a building block in integrated quantum optical architectures for computing and simulation where entanglement measurement and control plays an essential role.
@@
Figure 2: Sequential optimization methods for entanglement analysis in self-configuring optics. a-d. Sequential power optimization. a. Power optimization sequence: from $1$ to $N_{A}$ (on outputs of network $A$), then from $1$ to $N_{B}$ (on outputs of network $B$). b. Random matrix input example. c. Resulting eigenvector reconstruction from power optimization. d. Resulting simulated square roots of power outputs (meas. power) and singular values (sing. values) at the outputs of network $A$ (red) and $B$ (blue). e-h. Sequential coincidence counts optimization. e. Coincidence counts optimization sequence: from $1$ to $\text{min}(N_{A},N_{B})$ (on pairs of outputs ports from networks $A$ and $B$). f. Random matrix input example. g. Resulting eigenvector reconstruction from coincidence optimization. h. Resulting simulated square roots of coincidence counts (meas. coinc.) and singular values (sing. values). i. Square root of coincidence over iteration number, mapping to their corresponding Schmidt values. j. Fidelity of Schmidt modes in $A$ (red) and in $B$ (blue) over iteration number.
@@
In our numerical simulations, optimization is performed with a modified version of stochastic gradient descent kingma2014adam and the performance of the algorithm in shown in Figure 2(i,j) for the sequential coincidence optimization. Each sequential coincidence (from $1$ to $\text{min}(N_{A},N_{B})$) quickly converges to the corresponding singular value. The fidelity increases over each sequential optimization step, and reaches values $\approx 1$ in our numerical experiments (for both Schmidt vectors in $A$ and $B$, see fidelity definition in the SM).
We now discuss the computational cost (both in terms of physical resources and number of measurements) for the BSCN to perform Schmidt decomposition (e.g., to find the Schmidt mode corresponding to the largest Schmidt value). Importantly, since both methods rely on Rayleigh quotient optimization, their convergence can be formally proven, as we show in the SM, Section S7. This variational formulation of entanglement modal analysis also lends itself to stochastic gradient descent methods that have advantageous convergence speeds (which can be essentially independent of the state’s dimensionality in some implementations du2017gradient ). More generally, the number of measurements to achieve a given accuracy for a pair of Schmidt modes will scale like $O(N)$ (where $N$ is the Hilbert space dimension), and we show this linear scaling in simulations in the SI, Section S7. Additionally, our method naturally orders the $r$ Schmidt values in decreasing order, such that only $O(rN)$ physical resources (number of MZIs) and $O(rN)$ measurements are in principle required, which can be a substantial improvement for high-dimensional systems with low Schmidt numbers, as compared to the worst-case $O(N^{2})$ requirement using traditional Schmidt decomposition methods.
If single photons in either subspace are lost (e.g., absorbed or scattered), the input and output can mix with a vacuum state, which might result in errors in the power or coincidence measurements. In the SM, Section S3 we model such detection and insertion losses as coupling each output and input port to a reservoir. We find that output losses do not affect the optimization process, and can be readily calibrated out by measuring the transmission matrix of the network (as is typically done in entanglement distribution experiments lib2022quantum ). Non-uniform insertion losses at the input can distort the input state miller2013establishing ; roques2024measuring . In that case, the coincidence optimization can still learn the correct SVD of the distorted input state traversing the network, while the sequential power optimization does not (nonetheless, in both cases there is enough information to reconstruct the original state, if the transmission matrix is known). This problem can be alleviated by equalizing insertion losses, preventing distortions, and allowing both the power and coincidence optimizations to learn the SVD of the input quantum state. Large losses may however result in reducing the number of counts on the detector and slow down the training process. Other MZI mesh architectures may also alleviate the influence of losses and imperfections burgwal2017using ; miller2015perfect . This robustness to single photon losses provides further motivation to investigate the performance of our method in experimentally-relevant settings.
@@
Figure 4: Schmidt mode reconstruction and entangled light processing with bipartite self-configuring optics. a. Left: First four Schmidt modes reconstructed from BSCN parameters in the SPDC example of Figure 3d,g.. Right: BSCN architecture for reference, where each self-configuring layer is highlighted in a different color. The self-configuring layers used to reconstruct the Schmidt modes on the left are represented by the number and color of gray squares as insets. b. Once trained, BSCN can be used to generate separable light with a given modal shape by blocking all output ports except the two ports corresponding to that mode index (in networks $A$ and $B$) and feeding them as input to networks applying the inverse transformation. c. If trained on a highly degenerate input state (such that two singular values are approximately equal), BSCN can also be used to generate supermode Bell states (by blocking all output ports except the four ports corresponding to degenerate singular values in each network). An additional phase shifter can be added on one of the output ports to control the relative phase between the two terms of the pair.
Figure 5: Entanglement distribution through scattering media with BSCN. Proposed setup for quantum teleportation through scattering media with BSCN. A source $C$ is generating Bell states sent to Alice ($A$) and Bob ($B$) who analyze their single photon with their own SCN. A classical communication channel between the two allows them to perform coincidence measurements.
@@
In the absence of insertion and detection losses, and spectral mixing by the gratings and frequency shifters (as shown in SM Section S3, and discussed above, the BSCN is robust to such imperfections), the resulting input state to the BSCN is:
@@
A distinctive feature of our method is that the Schmidt modes (including their phases, see SM, Section S2C) are directly encoded into the parameters of the network after it has converged. Specifically, the $k$-th Schmidt mode is encoded in the network parameters (e.g., the phase shifts for various waveguide segments of MZIs) of layers $1,\ldots,k$ of both networks $A$ and $B$. We can therefore reconstruct the Schmidt modes (of the SPDC example from Figure 3(d)) as shown in Figure 4(a). The self-configuring layers used for the reconstruction of each mode are indicated with dashed square boxes in each panel.
@@
Entanglement distribution through scattering media with BSCN
By shaping the modal composition of entangled states, our approach can also be applied to quantum communication protocols. One famous example is that of quantum teleportation bouwmeester1997experimental between two agents Alice ($A$) and Bob ($B$), where a Bell state is provided by a third-party source $C$ (see Figure 5). A successful implementation of quantum teleportation entails the distribution of entanglement between the two parties $A$ and $B$, or, in other words, communication of each photon of the Bell state to each of the two parties. In real-life settings, this communication may experience scattering and mode mixing through different communication channels (e.g., due to random scattering). The scattering of quantum light through complex media is an active field of research lib2022quantum , with immediate applications to quantum communication and entanglement distribution protocols through scattering channels such as turbulent atmosphere and multimode fibers. Conventional schemes compensate for the scattering through adaptive wavefront shaping of the photon pairs by characterizing the channel using an auxiliary laser at or near the same photon frequency defienne2016two ; defienne2018adaptive ; wolterink2016programmable ; cao2020long or even the pump laser used to generate the photons lib2020real .
The BSCN offers an alternative approach based on automatic Schmidt decomposition via sequential coincidence optimization. Specifically, if both Alice and Bob possess a SCN that they use to process the photon they receive, and perform the Schmidt learning method described in this work, they can learn the basis in which Alice should perform measurement to teleport her state to Bob. To do so, a classical communication channel is necessary between Alice and Bob, but there is no need to perform any characterizations of the communication channels between the source to either Alice or Bob. More generally, the BSCN allows us to undo the influence of random scattering when distributing multimode entangled states. A formal derivation of this communication protocol is provided in the SM, Section S8.
In fact, in this protocol the BSCN finds the effective communication modes miller2013establishing ; seyedinnavadeh_determining_2023 between Alice and Bob. This idea can therefore be generalized to the Klyshko configuration shekel2024shaping , which replaces one of the photon detectors (say, Alice’s) by a classical input laser beam to characterize the effective scattering channel between Alice, the source, and Bob. Characterizing the communication channel in this manner to configure the BSCN could boost the convergence speed and detection bandwidth of the protocol, which can be critical for entanglement distribution through rapidly fluctuating media.
@@ Discussion
Our methods could be applied to entanglement analysis in other quantum systems, such as temporal entanglement in atom-photon coupling bogdanov2006schmidt and macroscopic quantum states of light, such as bright squeezed vacuum sharapova2015schmidt ; sharapova2018bright . In both cases, multimodal entanglement over temporal, spatial, or spectral degrees of freedom can be described on the Schmidt basis which reveals the fundamental mechanisms of entanglement generation. Many of these systems, including photon pairs from SPDC, rely on several control parameters of light or matter, which can be dynamically optimized to shape or maximize entanglement. Since the BSCN automatically learns the Schmidt decomposition of the input state, it could also serve as a form of feedback on the system’s control parameters to learn their influence on the entanglement structure of the input state. In particular, well established techniques for shaping photon pairs u2006generation ; dosseva2016shaping ; meyer2017limits ; graffitti2020direct ; morrison2022frequency ; hurvitz2023frequency ; shukhin2024two are often sensitive to experimental parameters such as beam alignment, poling fabrication, and even temperature. Our proposal could thus allow for robustness and tunability even in the presence of experimental imperfections.
The automatic learning of the input state’s Schmidt decomposition requires a pure input state of the network |$\ket{\psi_{\text{in}}}$. Several sources of noise and imperfections may result in impurity of the quantum state. First, stochastic losses through the network yield a mixture of pure states with different photon numbers. With the above-mentioned loss model, we show that post-selection on a given number state (e.g., by performing coincidence or single photon counts) mitigates the influence of this impurity on the performance of our algorithm (see SM, Section S3). Second, the input state might be mixed with other (random) states, resulting in a statistical mixture. In general, this would result in the emergence of cross-coincidence counts between output ports |$\ket{x_{j}}_{A}$ and |$\ket{x_{k}}_{B}$ where $j\neq k$ after diagonalization of the reduced density matrices performed by the BSCN (see SM, Section S6). Practically, this means that the BSCN might also be used to “detect” impurities in the input quantum state in this more general setting.
Our proposal provides a high-dimensional (two-qudit) photonic-hardware implementation of variational quantum singular value decomposition circuits bravo2020quantum ; wang2021variational which decompose bipartite entanglement of general multiple-qubit states. As a future direction, it will be interesting to generalize our architecture to bipartite entangled states of more than two photons, allowing for multiple-qudit variational quantum singular value decomposition circuits.
@@
I Ethics approval and consent to participate
Not applicable.
II Availability of data and materials
III Competing interests
@@
IV Funding
C. R.-C. is supported by a Stanford Science Fellowship. A.K. is supported by the VATAT-Quantum fellowship by the Israel Council for Higher Education; the Urbanek-Chodorow postdoctoral fellowship by the Department of Applied Physics at Stanford University; the Zuckerman STEM leadership postdoctoral program; and the Viterbi fellowship by the Technion. S. F. and D.A.B.M. acknowledge support by the Air Force Office of Scientific Research (AFOSR, grant FA9550-21-1-0312). D.A.B.M. also acknowledges support by the Air Force Office of Scientific Research (AFOSR, grant FA9550-23-1-0307).
V Authors’ contributions
C.R.-C., A.K., D.A.B.M., and S.F. conceived the idea. C.R.-C. and A.K. performed the numerical experiments. C.R.-C. and A.K. wrote the manuscript with inputs from all authors. S.F. and D.A.B.M. supervised the research.
VI Acknowledgements
The authors would like to thank Paul-Alexis Mor, Carson Valdez, Annie Kroo, Anna Miller, and Olav Solgaard for stimulating conversations.